# Nelson’s Stochastic Mechanics: Measurement, Nonlocality, and the Classical Limit

### 03 Apr 2026 v1

@@ Abstract
Nelson’s stochastic mechanics may be understood as a stochastic underpinning, or reconstruction, of nonrelativistic quantum mechanics, once the diffusion scale is fixed by $\hbar$ and the admissible states are restricted by the usual single-valuedness condition on the wavefunction. In this note I briefly indicate what this route achieves and why it remains conceptually attractive. Three advantages are emphasized. First, it supplies a clear configuration-space stochastic picture of the underlying processes. Second, it offers a markedly different perspective on measurement and nonlocality: in particular, collapse need not be treated as an extra axiom, and the nonlocality associated with entangled states is softened relative to the deterministic Bohmian guidance picture. Third, by tying quantumness to a diffusion scale, it naturally suggests a continuum of physical descriptions ranging from the strictly classical to the strictly quantum-mechanical regime. I conclude by proposing a natural distance scale in stochastic mechanics and examining its implications for testing possible limits of Bell correlations.
@@ 2 What Nelson’s route achieves
This is the first point that should be stated clearly: Nelson’s approach is best regarded as a stochastic underpinning or reconstruction of nonrelativistic quantum mechanics, not as a rival set of empirical predictions.

### 08 Apr 2026 v2

@@ Abstract
Nelson’s stochastic mechanics may be understood as a stochastic underpinning, or reconstruction, of nonrelativistic quantum mechanics, once the diffusion scale is fixed by $\hbar$ and the admissible states are restricted by the usual single-valuedness condition on the wavefunction. In this note I briefly indicate what this route achieves and why it remains conceptually attractive. Four advantages are emphasized. First, it supplies a clear configuration-space stochastic picture of the underlying processes. Second, the Born rule is built in from the outset, with $|\psi|^{2}$ arising as the probability density $\rho$ of the underlying diffusion process rather than as an independent postulate. Third, it offers a markedly different perspective on measurement and nonlocality: in particular, collapse need not be treated as an extra axiom, and the nonlocality associated with entangled states is softened relative to the deterministic Bohmian guidance picture. Fourth, by tying quantumness to a diffusion scale, it naturally suggests a continuum of physical descriptions ranging from the strictly classical to the strictly quantum-mechanical regime. I conclude by proposing a natural distance scale in stochastic mechanics and examining its implications for testing possible limits of Bell correlations.
@@ 2 What Nelson’s route achieves
An important advantage of Nelson’s stochastic mechanics is that the Born rule is not an independent postulate. The theory begins with a diffusion process $X(t)$ in configuration space, whose probability density is $\rho(x,t)$. When the Schrödinger equation is recovered within this framework, one finds (see eqn (2))
$\rho(x,t)=|\psi(x,t)|^{2}.$
The Born rule is therefore built into the theory at the foundational level: it expresses the distribution of the underlying stochastic process itself. This is conceptually preferable to the orthodox formulation, in which the rule $|\psi|^{2}$ must be added separately as a probabilistic postulate.
Thus, whereas orthodox quantum mechanics takes as foundational postulates the Schrödinger equation, the Born rule, and the single-valuedness of the wave function, Nelson’s stochastic mechanics improves on it in two respects. First, the Schrödinger equation is not postulated but reconstructed from an underlying stochastic dynamics. Second, the Born rule is not an independent axiom, since $|\psi|^{2}$ arises directly as the probability density of the diffusion process. Only one supplementary condition remains to be imposed, namely the usual single-valuedness of the wave function. This is the first point that should be stated clearly: Nelson’s approach provides a stochastic underpinning, or reconstruction, of orthodox quantum mechanics while retaining the same empirical predictions.